OpenStudy (anonymous):

In the equation y^2=4px find the value of p under each of the follow conditions; a) the distance from the vertex to the focus is 3 b) the focus is the point (-5, 0)

OpenStudy (anonymous):

for part b) i think p=-5 am i correct?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

i never really worked with that form, what does p represent?

OpenStudy (anonymous):

the focus point i think like example y^2=12x so 4p=12 p=12/4 p=3 the focus (3,0) directrix x=-3

OpenStudy (amistre64):

spose f(0,p) and d: y=-p (x-0)^2 + (y-p)^2 = (x-x)^2 + (y+p)^2 x^2 + y^2 -2py +p^2 = y^2 +2py + p^2 x^2 -2py = 2py x^2 = 4p y ok, p represents the distance from vertex to focus/directrix

OpenStudy (amistre64):

find the value of p, well if the distance given is 3 ... then p=3 :) and -5 seems fine for the other one yes

OpenStudy (anonymous):

ok so p is the distance

OpenStudy (amistre64):

part a can have 2 results. 3 and -3 since distance is not dependant on direction

OpenStudy (amistre64):

the equation, let p represent the distance AND direction from vertex to focus

OpenStudy (anonymous):

ok would u say that the latus rectum when divided by 2 gives the focal radius?

OpenStudy (amistre64):

the latrec being the line thru the focus roght?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):